# Recovery speed

### v1

@@ Abstract
This research proposes methods for formulating and guaranteeing the resilience of long short-term memory (LSTM) networks, which can serve as a key technology in AI system quality assurance. We introduce a novel methodology applying incremental input-to-state stability ($\delta$ISS) to mathematically define and evaluate the resilience of LSTM against input perturbations. Key achievements include the development of a data-independent evaluation method and the demonstration of resilience control through adjustments to training parameters. This research presents concrete solutions to AI quality assurance from a control theory perspective, which can advance AI applications in control systems.
marginparsep has been altered.
topmargin has been altered.
marginparpush has been altered.
The page layout violates the ICML style. Please do not change the page layout, or include packages like geometry, savetrees, or fullpage, which change it for you. We’re not able to reliably undo arbitrary changes to the style. Please remove the offending package(s), or layout-changing commands and try again.
Enhancing AI System Resiliency: Formulation and Guarantee
for LSTM Resilience Based on Control Theory
Sota Yoshihara* 1  Ryousuke Yamamoto* 2  Hiroyuki Kusumoto* 1  Masanari Shimura1
footnotetext: *Equal contribution 1Graduate School of Mathematics, Nagoya University, Nagoya-shi, Aichi, Japan 2Aisin Software, Kariya-shi, Aichi, Japan. Correspondence to: Sota Yoshihara <sota.yoshihara.e6@math.nagoya-u.ac.jp>.
Preprint.
@@ 1.1 Background
In recent years, artificial intelligence (AI) has become increasingly integrated into various aspects of our daily lives, from healthcare to autonomous vehicles. This widespread adoption has raised concerns regarding AI system reliability and the need for robust quality assurance measures. In response to these trends, multiple organizations and regulatory bodies have published guidelines for AI quality assurance. For instance, the European Union’s AI Act  European Parliament and of the Council (2024) and AI HLEG guidelines  High-Level Expert Group on Artificial Intelligence (2019) provide frameworks for ensuring AI safety and reliability. Although these guidelines have contributed to the gradual establishment of AI quality assurance processes, the development of specific evaluation metrics and implementation methodologies remains in its nascent stages.
@@
The development of these metrics and implementation methodologies is a critical issue, especially for control systems incorporating AI. Such systems interact directly with the physical world, potentially impacting human lives in extreme scenarios. These systems often employ time-series models such as recurrent neural networks (RNNs) and long short-term memory (LSTM) networks. These models can extract non-linear features from time-series data, store them as internal states, and utilize them to make inferences at each time step. Especially, LSTM surpasses other time series models in the accuracy of prediction Gers et al. (2002); Ma et al. (2015); Zhao et al. (2017). This ability is well-suited for predicting future outputs of nonlinear systems in model predictive control (MPC), and its capacity to track target values surpasses that of traditional control methods Wu et al. (2019); Chen et al. (2020); Kang et al. (2023). However, the state-space nature of these time-series models also presents significant risks. If a model receives temporarily anomalous input values, it may retain inaccurate state memories, potentially leading to persistent error in prediction that could significantly compromise the system’s safety and reliability. For theoretical and Experimental approaches in assurance of LSTM stability and quality, see also Section 2.2 and Section 2.3.
@@
The concerns arising from the use of time-series models for control systems can be mitigated by ensuring stability, robustness, and resilience in these models Dawson et al. (2022). Stability refers to consistent performance across similar inputs, which forms the foundation for robustness and resilience. Robustness refers to patience against perturbations, while resilience refers to the ability to return to a normal state after transitioning to an abnormal mode due to anomalous inputs. See also Section 2.1 for related work on these concepts. Time-series models with these guaranteed properties should exhibit consistent behavior under normal conditions. However, these properties cannot be adequately evaluated using conventional metrics focused on inference accuracy, such as precision, recall, and F1-score Sengupta & Friston (2018). Therefore, new mathematical formulations of evaluation metrics and techniques to satisfy these metrics are necessary. Although this issue might be addressed by comprehensively collecting and learning from data across various scenarios for evaluation, this approach is infeasible due to the difficulty of defining comprehensiveness and collecting data at scale.
Given this context, our research examines the mathematical formulation of the resilience of LSTM, which is different from data-driven approaches.
1.2 Methods
Figure 1: Schematic diagram of recovery time. The blue line depicts the system’s response to normal input without perturbations, while the orange and gray lines illustrate responses of systems with and without resilience to anomalous inputs subjected to instantaneous noise, respectively. We define the recovery time as the interval required for a system’s response to revert to its nominal behavior following a perturbation induced by noise. A resilient system is a system that has a finite recovery time.
@@ 1.2 Methods
  • To Introduce a novel mathematical formulation of the resilience, recovery time, defined as a duration required for a system to return to its normal state after transitioning to an abnormal mode due to anomalous inputs (See Figure 1).
@@
The key tool to achieve our goals is the incremental input-to-state stability ($\delta$ISS) Angeli (2002); Bayer et al. (2013), because $\delta$ISS guarantees the resilience of LSTM. Specifically, even if temporary perturbations in the input sequence disrupt inference, the state variables can converge to the same inference values that would have been obtained without perturbations. Sufficient conditions for the LSTM to exhibit the $\delta$ISS properties have already been derived as inequalities of parameters Terzi et al. (2021); Bonassi et al. (2023). However, it is not clarified in these studies how long it takes for state variables to converge. We will refine and improve upon these previous studies to achieve the above three goals.
@@ 1.3 Contribution
The main contributions of this study are as follows.
Theoretical Contributions:
  • Refinement of an evaluation of LSTM’s invariant sets which form the foundation for stability analysis (Section 4).
@@ 2.1 Control System and Stability Theory
In control theory, stability and resilience are important properties. Furthermore, it is well-established that the stability of controlled objects fundamentally determines the overall stability and resilience of control systems. For instance, systems possessing input-to-state stability (ISS) properties achieve global asymptotic stability under closed-loop control Jiang & Wang (2001). Moreover, systems with $\delta$ISS characteristics acquire robustness when implemented with model predictive control (MPC) Bayer et al. (2013). Therefore, in control methods such as MPC where controlled objects are modeled as dynamical systems, stability requirements for the dynamical systems are essential. While stability can be comprehensively guaranteed by general theory for linear systems, practical applications require specific and detailed engineering for a nonlinear system. This research focuses on LSTM networks as a specific case of nonlinear systems.
@@ 2.2 Theoretical LSTM Stability Assurance
Theoretical approaches have investigated sufficient conditions for LSTM networks to satisfy various stability properties. For single-layer LSTMs, conditions for global asymptotic stability Deka et al. (2019), ISS Bonassi et al. (2020), and $\delta$ISS Terzi et al. (2021) have been successively established. Research on sufficient conditions for $\delta$ISS has been extended to multilayer LSTM architectures Bonassi et al. (2023).
However, existing studies face two significant limitations: while they can guarantee recovery, they cannot provide estimates of recovery time, and their guaranteed operational ranges are relatively narrow. Our research addresses these challenges by advancing the stability theory of LSTM to enable both quantitative evaluation and adjustment of LSTM resilience. Furthermore, we aim to expand the guaranteed operational range of these stability conditions.
@@ 2.3 Experimental LSTM Quality Assurance
This paper discusses theoretical LSTM quality assurance; however, there are also some experimental methods. For example,  Wehbe et al. (2018) evaluated the robustness of LSTM inference using noise-added datasets.  Ahmed et al. (2020) proposed creating robust LSTMs by expanding training datasets to include abnormal scenes. Grande et al. (2021) proposed a data-driven stability assurance method.
These data-driven verification methods, like the ones mentioned above, are comprehensive when sufficient data are available, yet face challenges in determining adequate data quantities and provide only sparse guarantees for nonlinear models like LSTMs. While theoretical verification offers limited but exhaustive guarantees within its scope, these approaches are complementary rather than contradictory. Optimal verification should utilize both theoretical guarantees as possible and data-driven methods elsewhere.
@@ 3 Preliminary
In this section, we introduce essential mathematical preliminaries, $\delta$ISS, which is an important stability concept, and the LSTM.
@@ 3.2 Stability
Definition 3.2 ($\delta$ISS, Angeli 2002; Bayer et al. 2013).
@@ 3.3 LSTM Network
$\quad\odot\phi(c^{(l)}(t+1)),$ (4)
$y(t)$ $=U_{y}h^{(L)}(t)+b_{y},$ (5)
$x^{(l)}(t)$ =$=\begin{cases}x&\text{for }l=1,\\ h^{(l-1)}(t+1)&\text{for }l=2,\ldots L,\end{cases}$ (6)
@@ 4 $\delta$ISS of LSTM
In this section, we relax the existing $\delta$ISS condition of the LSTM (Terzi et al., 2021) by refinement the invariant set of the LSTM dynamics. This relaxation contributes to an accurate evaluation of the recovery time introduced in Section 5.
@@ 4.1 Invariant Set in LSTM
where $\bm{1}_{n}$ is a $n$-dimensional vector of all components being one, and $v_{+}$ denotes a vector $(\max\{v_{(i)},0\})_{i\in[n_{c}]}$ for a vector $v$. Note that $x_{\mathrm{max}}=1$ if $n\geq 2$. In addition, we define the sequences $\{\overline{\sigma}^{(l)}_{*}(k)\}_{k=0}^{\infty}$, $\{\overline{\phi}^{(l)}_{c}(k)\}_{k=0}^{\infty}$, $\{\eta^{(l)}(k)\}_{k=-1}^{\infty}$, and $\{\overline{c}^{(l)}(k)\}_{k=0}^{\infty}$ using following recursive formulas
@@
The proof is given in Appendix  B.1. We note that the invariant set $\mathcal{S}^{(l)}(0)$ is considered in the literatures (Terzi et al., 2021; Bonassi et al., 2023).
@@ 4.2 $\delta$ISS of LSTM
Let $L\in\mathbb{N}$ and $k\in\mathbb{Z}_{\geq 0}$. We assume that $s^{(l)}(0)\in\mathcal{S}^{(l)}(k)$ for all $l\in[L]$. The LSTM is $\delta$ISS if $\rho(A^{{(l)}}_{s}(k))<1$ for all $l\in[L]$, where,
@@
The proof is given in Appendix B.3. We note that a similar relaxation can be obtained in ISS (see Appendix B.3).
@@
    Theorem 4.2 coincides with existing results (Terzi et al., 2021; Bonassi et al., 2023) for the case $k=0$. Monotonically decreasing of $\rho(A^{(l)}_{s}(k))$ makes it possible to relax the constraints on the model in the case of $k>0$.
@@ 5 Recovery Time
This section introduces a novel formulation and an evaluation methodology for the recovery time of LSTMs.
We begin by proposing a rigorous mathematical formulation that defines the recovery time of LSTMs in Section 5.1. This definition accurately expresses the time required for the system to recover from disturbances and return to its normal state. However, the data dependence of this definition poses challenges for evaluation in practical applications.
To overcome this limitation, we derive a practical and computable upper bound for the recovery time in Section 5.2. This upper bound constitutes the primary novel contribution of our research, offering a fresh perspective on LSTM performance evaluation.
@@ 5.1 Definition
$=\min\{t_{1}\geq t_{0}:|y(t,s(0),x(0:t))$ (19)
$\hskip 40.0pt-y(t,s(0),\widehat{x}(0:t))|\leq e\ \text{for any}\ t\geq t_{1}\}-t_{0},$ (20)
@@
Evaluating the recovery time $T_{R}$ is difficult because it depends on the input data. In this paper, we utilize the function $\beta$ of $\delta$ISS in Definition 3.2.
@@ 5.2 Upper Bound of Recovery Time
where $U$ is a suitable unitary matrix (note that any matrix is triangulizable by some regular matrix $U$). Also, we regard the term of $\sum_{l=1}^{0}$ as 00.
This representation is given in the proof of Theorem 4.2 (see (167) and (130) in Appendix B.3).
@@
$\leq\min\{t_{1}\geq 0\ :\ \tilde{\beta}(2\bar{s}(k),t;k)\leq e/\|U_{y}\|\ \text{for any}\ t\geq t_{1}\}$ (36)
where $\bar{s}(k)=(\bar{s}^{(1)}(k),\dots,\bar{s}^{(L)}(k))$ and $\bar{s}^{(l)}(k)=\|(\bar{c}^{(l)}(k),\phi(\bar{c}^{(l)}(k))\bar{\sigma}^{(l)}_{o}(k))\|$
This is obvious from Definition 5.1 and Proposition 5.2.
@@ 6 Proposal
The implementation framework for the machine learning algorithm proposed in this study provides a comprehensive strategy for translating theoretical foundations into practical applications.
@@ 6.1 Estimate
Aim: Quantitatively assess the resilience of LSTMs.
Method: We propose to evaluate the obtained model’s resilience by the following index:
@@
Use case: We propose to establish this metric as a formal requirement for assessing LSTM quality, contextualized within the framework of control periods.
@@ 6.2 Training
Aim: Build an LSTM with a specified degree of resilience.
Method: We propose to control the model’s resilience (i.e., $T_{R}$) during training by the following loss function:
$Loss$ $=\mathcal{L}+\lambda\Phi(\theta,\varepsilon)$ (37)
@@
where, $\mathcal{L}$ denotes the existing loss function, $\Phi$ denotes the penalty term for the resilience, $\lambda>0$ denotes the penalty intensity, and $\varepsilon\in[0,0.5]$ denotes the control parameter for the resilience. The resilience of LSTM can be ensured through the design of the loss function, excluding modifications to the model architecture. In addition, the degree of resilience can be precisely calibrated by adjusting the parameter $\varepsilon$. In fact, as training progresses, $\Phi(\theta,\varepsilon)$ is expected to approach 0, and after training is completed, the model is expected to satisfy $\rho(A^{(l)}_{s}(k))\thickapprox 1-\varepsilon$. Based on Theorems 4.2 and 5.3, $T_{R}$ is expected to decrease as $\rho(A^{(l)}_{s}(k))$ decreases. Consequently, increasing the value of $\varepsilon$ enhances the resilience of the model. This amplification of resilience concurrently imposes stronger constraints on the model, potentially leading to a trade-off between resilience and model accuracy.
Use case: We tune $\varepsilon$ to balance resilience and accuracy requirements, observing the trade-off and selecting an optimal model.

### v4

# Enhancing AI System Resiliency: Formulation and Guarantee for LSTM Resilience Based on Control Theory

@@ Abstract
This paper proposes a novel theoretical framework for guaranteeing and evaluating the resilience of long short-term memory (LSTM) networks in control systems. We introduce recovery time as a new metric of resilience in order to quantify the time required for an LSTM to return to its normal state after anomalous inputs. By mathematically refining incremental input-to-state stability ($\delta$ISS) theory for LSTM, we derive a practical data-independent upper bound on recovery time. This upper bound gives us resilience-aware training. Experimental validation on simple models demonstrates the effectiveness of our resilience estimation and control methods, enhancing a foundation for rigorous quality assurance in safety-critical AI applications.
@@ 1.1 Background
In recent years, artificial intelligence (AI) has become increasingly integrated into various aspects of our daily lives, from healthcare to autonomous vehicles. This widespread adoption has raised concerns regarding AI system reliability and raised the need for robust quality assurance measures. In response to these trends, multiple organizations and regulatory bodies have published guidelines for ensuring AI safety and reliability, such as the European Union’s AI Act (European Parliament and of the Council,, 2024) and AI HLEG guidelines (High-Level Expert Group on Artificial Intelligence,, 2019). Although these guidelines have contributed to the gradual establishment of AI quality assurance processes, the development of specific evaluation metrics and implementation methodologies remains in its nascent stages. The development of these metrics and implementation methodologies is a critical issue, especially for control systems incorporating AI, because such systems interact directly with the physical world, potentially impacting human lives in extreme scenarios.
@@
In this paper, we focus on the control of time-series models, particularly long short-term memory (LSTM) networks, having real-time capability, which is an important property in control systems. Although typical time-series models can extract nonlinear features from time-series data, store them as internal states, and utilize them to make inferences at each time step, especially, LSTM surpasses other time-series models in the accuracy of prediction (see, e.g., Gers et al.,, 2002; Ma et al.,, 2015; Zhao et al.,, 2017). This ability is well-suited for predicting future outputs of nonlinear systems in model predictive control (MPC), and therefore LSTM’s capacity to track target values is better than traditional control methods (see, e.g., Wu et al.,, 2019; Chen et al.,, 2020; Kang et al.,, 2023). Moreover, LSTM can be implemented on Field Programmable Gate Array (FPGA), which are suitable for machine embedding (Guan et al.,, 2017). However, the state-space nature of time-series models also presents significant risks: If a model receives temporarily anomalous input values, it may retain inaccurate state memories, potentially leading to persistent errors in prediction that could significantly compromise the system’s safety and reliability.
@@
The concerns arising from the use of time-series models for control systems can be mitigated by ensuring stability, robustness, and resilience in these models (see, e.g., Dawson et al.,, 2022). Stability refers to consistent performance across similar inputs, which forms the foundation for robustness and resilience. Robustness refers to patience against perturbations, while resilience refers to the ability to return to a normal state after transitioning to an abnormal mode due to anomalous inputs. See also Section 2.1 for related work on these concepts. Time-series models with these properties should exhibit consistent behavior under normal conditions. However, these properties cannot be adequately evaluated using conventional metrics focused on inference accuracy, such as precision, recall, and F1-score (Sengupta and Friston,, 2018). Therefore, it is necessary to formulate new evaluation metrics and develop techniques to achieve the required metrics. This issue might be addressed by comprehensively collecting and learning from data across various scenarios for evaluation, but there are difficulties in defining comprehensiveness and collecting data at scale. Given this context, we mathematically formulate the resilience of LSTM, which is different from data-driven approaches. For theoretical and experimental approaches in assurance of LSTM stability and quality, see also Section 2.2 and Section 2.3.
1.2 Objective
@@
  • To introduce a novel mathematical formulation of the resilience, recovery time, defined as the duration required for a system to return to its normal state after transitioning to an abnormal mode due to anomalous inputs (see Figure LABEL:fig:image_recovery_time).
@@
The key tool to achieve our goals is the incremental input-to-state stability ($\delta$ISS) (Angeli,, 2002; Bayer et al.,, 2013), because $\delta$ISS guarantees the resilience of LSTM. Specifically, even if temporary perturbations in the input sequence disrupt inference, the state variables can converge to the same inference values that would have been obtained without perturbations. Sufficient conditions for the LSTM to exhibit the $\delta$ISS properties have already been derived as inequalities of parameters (Terzi et al.,, 2021; Bonassi et al.,, 2023). However, it is not clarified in these studies how long it takes for state variables to converge. We will refine and improve upon these previous studies to achieve the above three goals.
@@ 1.3 Contribution
The main contributions of this study are as follows:
Theoretical Contributions
  • Refinement of an evaluation of LSTM’s invariant sets, which form the foundation for stability analysis (Section 4).
@@ 2.1 Control System and Stability Theory
In control theory, stability and resilience are important properties. Furthermore, it is well-established that the stability of controlled objects fundamentally determines the overall stability and resilience of control systems. For instance, systems possessing input-to-state stability (ISS) properties achieve global asymptotic stability under closed-loop control (Jiang and Wang,, 2001). Moreover, systems with $\delta$ISS characteristics acquire robustness when implemented with model predictive control (MPC) (Bayer et al.,, 2013).
Therefore, in control methods such as MPC where controlled objects are modeled as dynamical systems, stability requirements for the dynamical systems are essential. While stability can be comprehensively guaranteed by general theory for linear systems, practical applications require specific and detailed engineering for a nonlinear system. This research focuses on LSTM networks as a specific case of nonlinear systems.
@@ 2.2 Theoretical LSTM Stability Assurance
Theoretical approaches have investigated sufficient conditions for LSTM networks to satisfy various stability properties. For single-layer LSTMs, conditions for global asymptotic stability (Deka et al.,, 2019), ISS (Bonassi et al.,, 2020), and $\delta$ISS (Terzi et al.,, 2021) have been successively established. Research on sufficient conditions for $\delta$ISS has been extended to multilayer LSTM architectures (Bonassi et al.,, 2023).
However, existing studies face a significant limitation: while they can guarantee recovery, they cannot provide estimates of time to recover. Our research addresses these challenges by advancing the stability theory of LSTM to enable both quantitative evaluation and adjustment of LSTM resilience.
@@ 2.3 Experimental LSTM Quality Assurance
This paper discusses theoretical LSTM quality assurance; however, there are also some experimental methods. For example,  Wehbe et al., (2018) evaluated the robustness of LSTM inference using noise-added datasets.  Ahmed et al., (2020) proposed creating robust LSTMs by expanding training datasets to include abnormal scenes. Grande et al., (2021) proposed a data-driven stability assurance method.
These data-driven verification methods, like the ones mentioned above, are comprehensive when sufficient data are available, yet face challenges in determining adequate data quantities and provide only sparse guarantees for nonlinear models like LSTMs. Theoretical verification offers limited but exhaustive guarantees within its scope, and is therefore complementary to, rather than in conflict with, a data-driven approach.
@@ 3 Preliminary
In this section, we introduce essential mathematical preliminaries, an important stability concept: $\delta$ISS, and the LSTM.
@@ 3.2 Stability
Definition 3.2 ($\delta$ISS, Angeli, 2002; Bayer et al., 2013).
@@ 3.3 LSTM Network
$\quad\odot\phi(c^{(l)}(t+1)),$ (3c)
$y(t)$ $=U_{y}h^{(L)}(t)+b_{y},$ (3d)
$x^{(l)}(t)$ =$=\begin{cases}x&\text{for }l=1,\\ h^{(l-1)}(t+1)&\text{for }l=2,\ldots L,\end{cases}$ (3e)
@@ 4 $\delta$ISS of LSTM
In this section, we relax the existing $\delta$ISS condition of the LSTM (Terzi et al.,, 2021) by refining the invariant set of the LSTM dynamics. This relaxation contributes to an accurate evaluation of the recovery time introduced in Section 5.
@@ 4.1 Invariant Set in LSTM
where $\bm{1}_{n}$ is a $n$-dimensional vector of all components being one, and $v_{+}$ denotes a vector $(\max\{v_{(i)},0\})_{i\in[n_{c}]}$ for a vector $v$. Note that $x_{\mathrm{max}}=1$ if $n\geq 2$. In addition, we define the sequences $\{\overline{\sigma}^{(l)}_{*}(k)\}_{k=0}^{\infty}$, $\{\overline{\phi}^{(l)}_{c}(k)\}_{k=0}^{\infty}$, $\{\eta^{(l)}(k)\}_{k=-1}^{\infty}$, and $\{\overline{c}^{(l)}(k)\}_{k=0}^{\infty}$ using following recursive formulas:
@@
Proof sketch.
Since the activation functions are strictly monotonically increasing and bounded, the sequences are monotonically decreasing and converge to some limit. This proposition is proved by induction on $k$. ∎
The exact proof is described in Appendix B.1. We note that the invariant set $\mathcal{S}^{(l)}(0)$ is considered in the literature (Terzi et al.,, 2021; Bonassi et al.,, 2023).
@@ 4.2 $\delta$ISS of LSTM
Proof sketch.
The matrix $A_{s}^{(l)}(k)$ is obtained by the invariant set in Proposition 4.1, Lipschitz continuity of the activation functions, triangle inequality, and norm properties. The monotonic decrease of $\rho(A^{{(l)}}_{s}(k))$ is demonstrated using the Perron–Frobenius theorem. ∎
The rigorous proof is described in Appendix B.3, and where we also see that a similar relaxation can be obtained for ISS.
@@
    Theorem 4.2 coincides with existing results (Terzi et al.,, 2021; Bonassi et al.,, 2023) for the case $k=0$. Monotonically decreasing of $\rho(A^{(l)}_{s}(k))$ makes it possible to relax the constraints on the model in the case of $k>0$.
@@ 5 Recovery Time
In this section, we define a novel mathematical formulation, the recovery time of LSTM, which accurately expresses the time required for the system to recover from disturbances and return to its normal state. This definition makes it difficult to evaluate in practical applications due to data dependence, but we overcome this limitation by deriving a practical and computable upper bound for the recovery time. This upper bound constitutes the primary novel contribution of our research, offering a fresh perspective on LSTM performance evaluation.
@@ 5.1 Definition
$=\min\{t\geq t_{0}:\|y(t^{\prime},s(0),x(0:t^{\prime}))$ (16)
$\hskip 40.0pt-y(t^{\prime},s(0),\widehat{x}(0:t^{\prime}))\|\leq e\ \text{for any}\ t^{\prime}\geq t\}-t_{0},$ (17)
@@
Evaluating the recovery time $T_{R}$ is difficult because it depends on the input data, then we utilize the function $\beta$ of $\delta$ISS in Definition 3.2.
@@ 5.2 Upper Bound of Recovery Time
where $U$ is a suitable unitary matrix (note that any matrix is triangulizable by some regular matrix $U$). Also, we regard the term of $\sum_{l=1}^{0}$ as 0.
Proof sketch.
This representation is given in the proof of Theorem 4.2, specifically by repeating the evaluation of the inputs of each layer with the state variables of front layers according to (3e). ∎
@@
$\leq\min\{t\geq 0:\tilde{\beta}(2\sqrt{n_{c}}\bar{s}(k),t^{\prime};k)\leq e/\|U_{y}\|\ \text{for any}\ t^{\prime}\geq t\},$ (33)
where $\bar{s}(k)=(\bar{s}^{(1)}(k),\dots,\bar{s}^{(L)}(k))$ and $\bar{s}^{(l)}(k)=\|(\bar{c}^{(l)}(k),\phi(\bar{c}^{(l)}(k))\bar{\sigma}^{(l)}_{o}(k))\|$.
Proof sketch.
This is from Definition 5.1, Proposition 5.2 (shifting the initial time by $t_{0}$), and the fact that the diameter of the invariant set is $2\sqrt{n_{c}}\bar{s}^{(l)}(k)$. ∎
Detailed proof is stated in Appendix B.6.
@@ 6 Proposal
We provide a comprehensive strategy for translating theoretical foundations into practical applications.
@@ 6.1 Estimate
Aim
Quantitatively assess the resilience of LSTMs.
Method
Evaluate the obtained model’s resilience using the following index:
@@
Use case
We propose to establish this metric as a formal requirement for assessing LSTM quality, contextualized within the framework of control periods.
@@ 6.2 Training
Aim
Build an LSTM with a specified degree of resilience.
Method
Control the model’s resilience (i.e., $T_{R}$) during training by the following loss function:
$Loss$ $=\mathcal{L}+\lambda\Phi(\theta,\varepsilon),$ (34)
@@
where $\mathcal{L}$ denotes the existing loss function, $\Phi$ denotes the penalty term for the resilience, $\lambda>0$ denotes the penalty intensity, and $\varepsilon\in[0,0.5]$ denotes the control parameter for the resilience. The resilience of LSTM can be ensured through the design of the loss function, without modifications to the model architecture. In addition, the degree of resilience can be precisely calibrated by adjusting the parameter $\varepsilon$. In fact, as training progresses, $\Phi(\theta,\varepsilon)$ is expected to approach 0, and after training is completed, the model is expected to satisfy $\rho(A^{(l)}_{s}(k))\thickapprox 1-\varepsilon$. Based on Theorems 4.2 and 5.3, $T_{R}$ is expected to decrease as $\rho(A^{(l)}_{s}(k))$ decreases. Consequently, increasing the value of $\varepsilon$ enhances the resilience of the model. This amplification of resilience concurrently imposes stronger constraints on the model, potentially leading to a trade-off between resilience and model accuracy.
Use case
We tune $\varepsilon$ to balance resilience and accuracy requirements, observing the trade-off and selecting an optimal model.